OpenStudy (anonymous):

Algebra 1: I can't figure this out and it's driving me insane; Have about 20 questions like it for my homework Solve the system of equations. y = –0.04x² + 5.2x y = 1.12x + 1.6

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

equate both the eq.

OpenStudy (anonymous):

in other words,substitute value of y from 1 eq. into other

OpenStudy (anonymous):

and solve

OpenStudy (anonymous):

I'm not understanding how to do this. I substituted and got this: 1.12x+1.6=-0.04x^2+5.2 but I don't know what to do next.

OpenStudy (cwrw238):

first simplify it notice that you have + 1.6 on left and + 5.2 on right side so subtract + 5.2 from both sides: 1.12x+1.6=-0.04x^2+5.2 1.12x + 1.6 - 5.2 = -0.04x + 5.2 - 5.2 so the 5.2 disappears from the LHS can you do the math on these 2 sides of the equation?

OpenStudy (cwrw238):

what is 1.6 - 5.2?

OpenStudy (anonymous):

Um.. -3.6?

OpenStudy (cwrw238):

right so now we have 1.12x - 3.6 = -0.04 x^2 ok?

OpenStudy (anonymous):

Okay. Then do I graph that?

OpenStudy (cwrw238):

you can solve it by graphing it is a Quadratics equation which is written in the form ax^2 + bx + c = 0 which can be solved algebraically rearranging we get 0.04x^2 + 1.12x - 3.6 = 0 are you familiar with these equations?

OpenStudy (anonymous):

Oh yes. I can solve it from here. Thank you!

OpenStudy (cwrw238):

good you can use the quadratic formula

OpenStudy (cwrw238):

it will probably give you 2 values for x ,then plugging in these values into one of the original equations will give you the values of y